Plankton Company Balance Sheet as of Dec 31, 2017 (in $)

Cash                             112,500            Accounts Payable                       150,000

Receivables                   336,000            Notes payable                              84,000

Inventories                    241,500            Other Current Liabilities          117,000

Total CA                      690,000             Total CL                                 351,000

                                                                 L-T debt                                   256,500

Net Fixed Assets           292,500             Common Equity                        375,000

Total Assets                  982,500      Total L+E               982,500

Plankton Company Income Statement (year ended Dec 31, 2017—in $)

Sales                                                                                                                 1,607,000

Cost of Goods Sold

            Materials                                  717,000

            Labor                                        453,000

            Heat, Light, and Power           68,000

            Indirect Labor                           113,000

            Depreciation                             41,500                                                1,392,000

Gross Profit                                                                                                       214,500

Selling Expenses                                                                                                  115,000

G&A Expenses                                                                                                      30,000

            EBIT                                                                                                          69,500

Interest Expense                                                                                                24,500

            EBT                                                                                                      45,000

Taxes (40%)                                                                                                       18,000

Net Income                                                                                                        27,000

Use the data above for the Plankton Company to calculate the following ratios:

The answers must be carried out to two decimal points.

Current ratio:

Quick ratio:

Inventory Turnover:

Total Assets Turnover:

Profit Margin:

ROA:

ROE:

ROIC:

Total Debt/Total Capital:

Using the DuPont Equation please calculate the ROE? Must show the calculation in order to get credit.
Answer #1

As per rules I am answering the first 4 subparts of the question

Current ratio= Current assets/ current liabilities

= 690000/351000

= 1.97

Quick ratio =(Cash + receivables)/current liabilities

= (112500+ 336000)/351000

=1.28

Inventory Turnover= COGS/ Inventory

= 1392000/241500

=5.76

Total Assets Turnover= Sales/ Total assets

= 1607000/982500

=1.64